A coin is tossed downward with an initial velocity of 20

meters per second at an altitude of 273

meters. Approximately how long does it take the coin to reach the​ ground? You will need to use the formula 4.9t2+v0t=s
Answer #1

coin is tossed downward with initialvel:20misec Altitude = 273m time boreach ground.formula 4-963 201.222 4:9ť? +206-273=02 t = -bt 94+204-273-0 496+2004-273030 6²qal za t=-200+1(20012 2x49 = -20 t :-1006143770 49 t = -100+14 3770 to-100-8193730 4It= -100+379.12 44 t = 5.697 Sec time to reach ground